# 3D Gaussian Splatting: Survey, Technologies, Challenges, and Opportunities

### v1

@@ 3.2 Photorealism
Other works attempt to reconstruct challenging scenes, such as reflective surfaces [37, 38, 39], and Relightable [40].
GaussianShader [37] reconstructs reflective surfaces by employing a hybrid color representation and integrating the specular GGX [41] and normal estimation module, which encompasses diffuse color, direct specular reflection, and a residual color component that accounts for phenomena such as scattering and indirect light reflections. Also, GaussianShader introduces shading attributes and a normal-geometry consistency constraint in 3DGS during the training process.
Mirror-3DGS [38] adds a learnable mirror attribute to determine the location of mirrors and introduces a Virtual Mirrored Viewpoint to aid in the reconstruction of mirror scenes based on the original 3DGS. And, Spec-Gaussian [39] replaces the original 3DGS with Anisotropic Spherical Gaussian to construct scenes with specular and anisotropic components and introduces an Anchor-based representation [26] for efficiency.
Relightable 3D Gaussian (R3DG) [40] represents 3D scenes using relightable points, each characterized by a normal direction, BRDF parameters, and incident lighting, where the incident lights are decomposed into global and local components with view-dependent visibilities. Then, a novel point-based ray-tracing technique, based on bounding volume hierarchy, is designed in R3DG for efficient visibility baking and real-time rendering with accurate shadow effects.
DeblurGS [42] tackles the challenge of inaccurate camera poses caused by severe blur, which impedes the effectiveness of Structure-from-Motion (SfM). DeblurGS optimizes sharp 3D scenes by estimating the 6-Degree-of-Freedom (6-DoF) camera motion for each blurry observation and synthesizing corresponding blurry renderings. It also introduces a Gaussian Densification Annealing strategy for training stability.
3.3 Generalization and Sparse Views
The challenges of generalization and reconstruction under sparse view settings have long attracted significant scholarly attention. Both implicit representations, such as NeRF, and explicit representations, like 3DGS, face substantial obstacles to practical application due to the necessity of re-training for each scene and the high demand for dense sample inputs. In this section, we discuss their experimental setups and provide a more detailed understanding of existing works.
3.3.1 Generalizable 3D Gaussian Splatting
The objective of existing generalizable 3D reconstruction or novel view synthesis tasks is to leverage extensive auxiliary datasets to learn scene-agnostic representations. In the research on NeRF [43, 44, 45], this process typically involves inputting a small number (1-10) of reference images with adjacent poses to infer the target image. The radiance field functions as an intermediary, effectively obviating the need for explicit scene reconstruction and transforming the task into a scene-agnostic novel view synthesis problem.
In contrast, the 3DGS’s explicit representation has led to a substantial body of works focused on using reference images to directly infer corresponding Gaussian primitives on a per-pixel basis, which are subsequently employed to render images from target views.
To achieve this, early works such as Splatter Image [46] propose a novel paradigm for converting images into Gaussian attribute images, thereby predicting the Gaussian primitives corresponding to each pixel. This paradigm is then extended to multiple reference images for better rendering performance.
However, unlike the generalization efforts in NeRF, the training difficulty for generalizable 3DGS increases significantly. Operations such as non-differentiable densification can negatively impact the generalization training process. Addressing these challenges, pixelSplat [47] aims to predict a probabilistic depth distribution from features extracted using an epipolar transformer structure [44] and sample this distribution to replace the non-differentiable components.
Moreover, methods based on Multi-View Stereo (MVS) have consistently achieved significant success in scene reconstruction and novel view synthesis, particularly with the introduction of cost volumes, which enhance the network’s spatial understanding. Similar to methods in MVSNeRF [48], MVSplat [49] proposes representing the cost volume using plane sweeps in 3D space and predicting the depths in sparse reference inputs, precisely locating the centers of Gaussian primitives. This approach provides valuable geometric cues for novel view synthesis.
Additionally, some studies [50, 51] focus on introducing triplane to achieve generalization capabilities.
The work [50] introduces a point-cloud representation decoder and a triplane representation decoder based on reference features. Through parallel decoding, they construct a hybrid representation that combines explicit point clouds with implicit triplane fields, enabling the Gaussian Decoder to directly predict Gaussian attributes after position query. Building on similar representations, AGG [51] introduces a pseudo-label-based initialization method and a multi-stage training strategy. This strategy encompasses coarse Gaussian generation followed by super-resolution using reference image features, thereby producing detailed outputs.
3.3.2 Sparse Views Setting
@@
Reconstructing from sparse inputs presents significant challenges, wherein the methodology of 3DGS is fundamentally analogous to that of NeRF, which aim to develop novel regularization strategies and integrate supplementary information, such as depth data.
Chung et al. [52] propose employing a monocular depth estimation model to predict depth maps, which are subsequently refined using SfM [53] for precise depth range. Additionally, their work incorporates depth smoothness loss and two modified techniques tailored for scenarios with limited samples. Building upon the depth supervision, FSGS [54] introduces a proximity-guided Gaussian upsampling method to increase the quantity and integrates new pseudo-views through a 2D prior model to further mitigate overfitting.
Subsequently, Touch-GS [55] extends this paradigm with tactile sensing in robotic perception applications. After aligning with monocular depth information, tactile sensing data effectively predict corresponding depth and uncertainty maps based on implicit surface representations, which are used to enhance the initialization and optimization processes.
Furthermore, DNGaussian [56] explores the problem from a regularization perspective, proposing two different regularization: hard depth and soft depth, to address the degradation of scene geometry. Then, DNGaussian introduces both global and local depth normalization methods to enhance sensitivity to subtle local depth variations.
Some studies have focused on the initialization and training strategy. GaussianObject [57] introduces an initialization strategy based on Visual Hull and an optimization method using distance statistical data to eliminate floaters. Additionally, GaussianObject designs a Gaussian repairing module, which includes data acquisition, training, and inference paradigms, thus leveraging pre-trained models to address occlusions and information loss under sparse viewpoint conditions.
4 Applications of 3D Gaussian Splatting
3DGS excels in various application domains due to its efficiency and photorealistic renderings, which include digital human reconstruction, Artificial Intelligence Generated Content (AIGC), and autonomous driving, among others. Building on prior explorations, 3DGS can be directly applied as a core technology in diverse research areas, effectively replacing traditional 3D representations.
4.1 Human Reconstruction
The application of 3DGS for digital human-related tasks, including human reconstruction, animation, and human generation, has garnered substantial attention within the research community. Recent works can be categorized according to the reconstructed parts.
4.1.1 Body Reconstruction
TABLE I: Comparison to Works for Body Reconstruction.
@@ 4.1.1 Body Reconstruction
    Pose-dependent Deformation     Novel Pose Animation     Fast Training     High Real-time Rendering(>60FPS)     Monocular Input     Super-resolution     Generalization
@@
Body reconstruction mainly focuses on reconstructing deformable human avatars from multi-view or monocular videos, as well as providing real-time rendering. We list comparisons of recent works in Tab. I.
@@
Most works [58, 60, 62, 61, 63] prefer to use well-preconstructed human models like SMPL [66] or SMPL-X [67] as strong prior knowledge. Nevertheless, SMPL is limited to introducing prior knowledge about the human body itself, thus posing challenges for the reconstruction and deformation of outward features such as garments and hair.
For the reconstruction of outward appearance, HUGS [60] utilizes SMPL and LBS only at the initial stage, allowing Gaussian primitives to deviate from the initial mesh to accurately represent garments and hair. Animatable Gaussians [65] uses a template that can fit outward appearance as guidance and leverages StyleGAN to learn the pose-dependent Gaussian maps, enhancing the capability for modeling detailed dynamic appearances. GaussianAvatar [61] employs a feature that captures the coarse global appearance, which is integrated with a pose feature. These combined features are then input into a decoder to predict parameters for Gaussian primitives. 3DGS-Avatar [63] introduces a non-rigid transformation module that utilizes a multi-level hash grid encoder to encode 3D position, and concatenates it with a pose latent to form the input for a shallow MLP network, which predicts offsets of Gaussian’s position, scale, and rotation under certain poses.
Furthermore, certain studies [60, 62] opt not to utilize this feature, thereby reducing the complexity of the deformation and facilitating faster rendering speeds. However, these approaches also introduce a degree of distortion.
Some studies project the problem space from 3D to 2D, thereby reducing complexity and enabling the utilization of well-established 2D networks for parameter learning.
ASH [64] proposes generating a motion-related template mesh via a deformation network and predicting a motion-related texture map from this mesh. Then, a 2D network predicts Gaussian parameters through the generated texture map. Similarly, Animatable Gaussians [65] projects a template mesh human model from canonical space onto two 2D planes, frontal and dorsal, learning Gaussian attributes in these spaces.
GPS-Gaussian [59] addresses the generalizable human novel view synthesis by introducing Gaussian Parameter Maps that can be directly regressed without per-subject optimizations. This approach is complemented by a depth estimation module that lifts 2D parameter maps into 3D space.
4.1.2 Head Reconstruction
In the field of human head reconstruction, like most pipelines using SMPL as a strong prior, the work by GaussianAvatars [68] integrates FLAME [69] meshes to provide prior knowledge with 3DGS to achieve superior rendering quality. This method compensates for the fine details and elements that the FLAME mesh fails to accurately depict or track. Nevertheless, Gaussian Head Avatar [70] critiques the use of FLAME meshes and Linear Blend Skinning (LBS) for facial deformation, noting that these relatively simple linear operations struggle to capture the nuances of intricate facial expressions. Instead, it proposes employing a MLP to directly predict the displacement of Gaussians as they transition from a neutral to a target expression. This approach facilitates the rendering of high-resolution head images, achieving resolutions up to 2K.
4.1.3 Others
@@
Additionally, 3DGS has introduced innovative solutions in other human-related areas. GaussianHair [71] focuses on the reconstruction of human hair, using linked cylindrical Gaussian modeling for the strands. At the same time, it specifically introduces the GaussianHair Scattering Model to further enhance the capture of structures, allowing the reconstructed strands to render with high fidelity under different lighting conditions.
The research presented in Gaussian Shadow Casting for Neural Characters [72] focuses on shadow computation under various perspectives and movements, which first reconstructs the volume of density, normals, and albedo values from input human poses and training images using NeRF. Subsequently, it fits the NeRF output density maps with a collection of anisotropic Gaussians. The Gaussian representation enables more efficient ray tracing and deferred rendering techniques, replacing traditional sampling processes and thus accelerating the computation of shadows.
Additionally, some research [73, 74] has explored the integration of 3DGS with generative models, which will be discussed in Sec. 4.2.
4.2 Artificial Intelligence-Generated Content (AIGC)
@@
Artificial Intelligence Generated Content (AIGC) leverages artificial intelligence technologies to autonomously produce content. Recently, there has been a surge in methodologies for generating 3D ($XYZ$) representations based on 3DGS. In this chapter, we systematically classify contemporary algorithms based on the types of prompts and the objects they generate. The categories include Image-to-3D Object Generation, Text-to-3D Object Generation, Multi-Object and Scene Generation, and 4D Generation ($XYZ-T$), as shown in Fig. 4. Below, we provide an overview of the relevant works within each of these categories.
@@
4.2.1 Text to 3D Objects
At present, a substantial body of research is dedicated to extending Score Distillation Sampling (SDS) [79], which plays a crucial role in this context, aiming to directly generate 3D representations with multi-view consistency using a distillation paradigm. To further elucidate SDS, we denote the 3D representations as $\theta$ and the differentiable rendering process as $g(\cdot)$, thereby representing the rendered image as $g(\theta)$. DREAMFUSION[79] ensures that the rendered images from each camera viewpoint adhere to the credible samples derived from the pre-trained diffusion model $\varphi$. In practice, they leverage the score estimation function $\epsilon_{\phi}(\bm{x_{t}},t,y)$ of the existing diffusion model, where $\epsilon_{\phi}$ predicts the sampled noise based on the noisy image $\bm{x_{t}}$ and the textual condition $y$. Therefore, the gradient of score distillation loss for $\theta$,
Subsequent works have widely adopted this, or improved variations thereof, as the primary supervision for generation.
Some works [75, 80, 81] focus on improving the framework to apply score distillation loss to 3DGS.
DreamGaussian [75], an early work integrating diffusion models with 3DGS, utilizes a two-stage training paradigm. Building upon Score Distillation Sampling (SDS), this approach ensures the geometric consistency of the generated models by extracting explicit Mesh representations from the 3DGS and refines texture in the UV space to enhance the quality of the renderings. The work by Chen et al. [80], conducted contemporaneously, introduces Point-E [82] (or other text-to-point-cloud models), along with 3D score distillation loss, to guide the 3D geometry generation in the first stage. And in the second stage, a density-based densification is employed to further refine the generation quality. Similarly, GaussianDreamer [81] adopts the same fundamental concept; however, it distinguishes itself by employing strategies such as noise point growth and color perturbation to address the issue of insufficient point cloud density during the initialization.
However, the mode-seeking paradigm of score distillation frequently leads to oversaturation, excessive smoothing, and lack-detail in the generated outcomes, which has been extensively discussed in related works on NeRF [83].
GaussianDiffusion [84] introduces variational Gaussian to mitigate the instability of 2D diffusion models and incorporates structured noise to enhance the 3D consistency.
There are also some works that focus on improving SDS.
LucidDreamer [85] addresses the challenges of over-smoothing and insufficient sampling steps inherent in traditional SDS. By introducing deterministic diffusion trajectories (DDIM [86]) and interval-based score matching mechanisms Eq. 9, it achieves superior generation quality and efficiency. Subsequently, Hyper-3DG [87] builds upon LucidDreamer [85] by introducing a hypergraph [88] to explore the relationships between patchified Gaussian primitives.
Similarly, LODS[89] analyzes the inherent inconsistency between the training and testing processes and the over-smoothing effect caused by large CFG (Classifier-Free Guidance) in SDS. To address these challenges, LODS proposes an additional set of learnable unconditional embeddings and low-rank parameters of LoRA (Low-Rank Adaptation) [90] for aligning distributions.
Score distillation loss can also be replaced in AIGC.
IM-3D [91] identifies optimization difficulties associated with score distillation loss. Thus, it attempts to fine-tune existing image-to-video generation models to enable the generation of multi-view spatially consistent images (videos). These generated multi-view images are then used as supervision for 3DGS generation. Similarly, LGM [92] proposes a novel paradigm to generate 3DGS from text or a single image. It generates multi-view images of the target using existing networks and reconstructs 3D models under different inputs using an asymmetric U-Net based architecture with cross-view self-attentions.
Works [93, 94] aim to generate using only feed-forward networks without the need for scene-specific training.
BrightDreamer [93] aims to incorporate generalizable representation, as described in 3.3.1, into text-to-3D generation, enabling the direct creation of 3D models without the need for retraining in specific scenes. BrightDreamer predicts positional offset after fixed initialization and introduces a text-guided triplane generator for the extracted textual features to predict other attributes of 3DGS, thereby achieving the conversion of arbitrary text into 3D models. GVGEN [94] focuses on a feed-forward setup without triplane, proposing GaussianVolumes as a structured lightweight representation for generalizable generation. Based on this representation, GVGEN generates a Gaussian Distance Field with a trained diffusion model and utilizes it to guide the prediction of corresponding attributes.
@@ 4.2.1 Text to 3D Objects
Some works [73, 74] also attempt to apply this generative paradigm to areas such as digital human generation.
HumanGaussian [74] combines RGB and depth rendering to improve the SDS, thereby jointly supervising the optimization of the structural perception of human appearance and geometry. Additionally, it introduces annealed negative prompt guidance and a scaling-based pruning strategy to address oversaturation and floating artifacts. In addition to the numerous works relying heavily on diffusion models, the work by Abdal et al. [73] proposes a novel paradigm that combines 3DGS with Shell Maps [95] and a 3D Generative Adversarial Networks (GANs) framework. By leveraging Gaussian Shell Maps, this approach rapidly represents human bodies and their corresponding deformations.
4.2.2 Image to 3D Object
Similar to works on NeRF, recent studies [96, 77] have also focused on generating entire 3DGS from a single image.
Following a process similar to DreamGaussian [75], Repaint123 [77] divides this procedure into coarse and fine optimization stages. In the coarse stage, it uses the pre-trained Zero-123 [97] as supervision and employs SDS to optimize a coarse 3DGS. During the fine stage, Repaint123 extracts the mesh representation from the first stage and proposes combining depth and reference images to guide the denoising process on novel view images, ensuring consistency across views. For overlapping and occluded regions between views, Repaint123 employs visibility-aware adaptive repainting methods to enhance repainting quality in these areas, which are then used for fine-tuning 3DGS.
FDGaussian [96] proposes a more straightforward approach by dividing the entire generation process into multi-view image generation and 3DGS reconstruction. During the generation phase, FDGaussian extracts 3D features from the image by decoupling orthogonal planes, optimizing the multi-view generation model based on Zero-1-to-3 [97]. In the reconstruction phase, it enhances the efficiency and performance of 3DGS through control optimization based on inter-Gaussian distances and a fusion strategy based on epipolar attentions.
4.2.3 Multi-Object and Scene Generation

### v2

@@ III-B Photorealism
Other works attempt to reconstruct challenging scenes, such as reflective surfaces [34, 35, 36, 37, 38, 39] and Deblurring [40, 41, 42, 43].
GaussianShader [34] reconstructs reflective surfaces by employing a hybrid color representation and integrating the specular GGX [44] and normal estimation module, which encompasses diffuse color, direct specular reflection, and a residual color component that accounts for phenomena such as scattering and indirect light reflections. The work [38] introduces a local Gaussian density mixture representation that combines geometry-guided ray interpolation with neural blending and end-to-end optimization to model high-frequency view-dependent reflections and refractions. Although these methods effectively manage complex reflective surfaces, they unavoidably sacrifice computational efficiency relative to vanilla 3DGS. Recently, the work [37] proposes a deferred shading method for 3DGS, which overcomes the challenge of normal estimation for environment map reflection and demonstrates enhanced efficiency by propagating accurate normals across neighboring Gaussians and per-pixel shading in screen space.
Existing deblurring approaches predominantly focus on motion blur or lens defocusing [40, 41, 42], introducing blur process modeling to enable sharp reconstruction supervised by blurred images. To address a wider range of scenarios, BAGS [43] introduces a Blur Agnostic robust modeling by a Blur Proposal Network and coarse-to-fine optimization scheme.
III-C Generalizable 3DGS
The objective of existing generalizable 3D reconstruction or novel view synthesis tasks is to leverage extensive auxiliary datasets to learn scene-agnostic representations.
The 3DGS’s explicit representation has led to a substantial body of works focused on using reference images to directly infer corresponding Gaussian primitives on a per-pixel basis, which are subsequently employed to render images from target views [45, 46]. To achieve this, early works such as Splatter Image [45] propose a novel paradigm for converting images into Gaussian attribute images. MVSplat [47] proposes representing the cost volume using plane sweeps in 3D space and predicting the depths in sparse reference inputs, precisely locating the centers of Gaussian primitives. While demonstrating generalization ability, this technique’s application is limited by its restricted synthesis range and the emergence of distractor-data. Subsequent works, FreeSplat [48] and DGGS [49], address these limitations through Pixel-wise Triplet Fusion strategy and distractor-free training and inference paradigms, respectively. Most recently, G3R [50] extends generalizable 3DGS to dynamic scenes, achieving generalized dynamic scene reconstruction through extra LiDAR data.
Furthermore, some studies [51, 52] focus on introducing triplane to achieve generalization capabilities , which infers Gaussian attributes by querying triplane features. Recent works [53, 54] seek to extend similar paradigms to large-scale 3D datasets, utilizing transformer-based architectures to enable direct 3D asset inference from sparse image inputs.
III-D Sparse Views Setting
@@
Reconstructing from sparse inputs presents significant challenges, wherein the methodology of 3DGS is fundamentally analogous to that of NeRF [55], which aim to develop novel regularization strategies and integrate supplementary information, such as depth data or Diffusion model [56, 57, 58, 59, 60, 61, 62, 63].
Specifically, the incorporation of depth data is crucial for 3DGS as an explicit representation, alleviating the model’s demand for spatial comprehension under sparse inputs. And multi-view trained diffusion models also can provide important prior knowledge for expanding sparse views into refined dense information. Nevertheless, these methods typically depend on the diffusion models’ ability to preserve 3D consistency. Additional regularization is commonly achieved through incorporating explicit spatial distribution rules or pseudo-views supervision. In addition to these common methods, some studies have focused on the initialization and optimization strategy. GaussianObject [64] introduces an initialization strategy based on Visual Hull and an optimization method using distance statistical data to eliminate floaters.
IV Applications of 3D Gaussian Splatting
IV-A Human Reconstruction
IV-A1 Body Reconstruction
TABLE II: Comparison to works for Body Reconstruction.
@@ IV-A1 Body Reconstruction
    Pose-dependent Deformation     Novel Pose Animation     Fast Training     Rendering > 60FPS     Monocular Input     Super-resolution     Generalization
@@
Body reconstruction mainly focuses on reconstructing deformable human avatars from multi-view or monocular videos [73], as well as providing real-time rendering. We list comparisons of recent works in Tab. II.
@@
Most works [65, 67, 69, 68, 70] prefer to use well-preconstructed human models like SMPL [74] or SMPL-X [75] as strong prior knowledge. Nevertheless, SMPL is limited to introducing prior knowledge about the human body itself, thus posing challenges for the reconstruction and deformation of outward features such as garments and hair. For the reconstruction of outward appearance, HUGS [67] utilizes SMPL and LBS only at the initial stage, allowing Gaussian primitives to deviate from the initial mesh to accurately represent garments and hair. Additionally, Animatable Gaussians [72] uses a template that can fit outward appearance as guidance and leverages StyleGAN to learn the pose-dependent Gaussian maps, enhancing the capability for modeling detailed dynamic appearances.
Some studies project the problem space from 3D to 2D, thereby reducing complexity and introducing well-established 2D networks for parameter learning [71, 72]. Among them, ASH [71] generates a motion-related template mesh via a deformation network and predicts Gaussian parameters through a 2D network using a motion-related texture map derived from this mesh. These methods are often limited by the difficulty of isolating clothing from the reconstructed 3DGS.
Although existing methods achieve high-fidelity reconstruction, their applications are limited by scene-specific training requirements, prompting research efforts toward generalizable human reconstruction [66].
IV-A2 Head Reconstruction
In the domain of human head reconstruction, the works [76, 77, 78, 79, 80], following the prevalent approach of utilizing SMPL as a strong prior, incorporates FLAME [81] meshes to provide geometric guidance or coarse reconstruction for 3DGS, achieving superior rendering quality. However, Gaussian Head Avatar [82] challenges the conventional use of FLAME meshes and LBS for facial deformation, arguing that these simplified linear operations are inadequate for capturing intricate facial expressions. As an alternative, it introduces an MLP-based approach that directly predicts Gaussian displacements during the transition from neutral to target expressions, enabling high-resolution head rendering at up to 2K resolution. And then, the works [83, 84] replace FLAME-based initialization with geometric guidance derived from signed distance field and DMTet. Several works [85, 86] focus on rendering efficiency, achieving frame rates exceeding 300 FPS. For downstream tasks, various studies [87, 88, 89] seek to integrate audio for controlling dynamic head reconstruction, thereby achieving audio-visual synchronization.
IV-A3 Others
@@
3DGS has introduced innovative solutions in other human-related areas [90, 91]. GaussianHair [90] focuses on the reconstruction of human hair, using linked cylindrical Gaussian modeling for the strands. Additionally, some research [92, 93, 94] have explored the integration of 3DGS with generative models.
IV-B Artificial Intelligence-Generated Content (AIGC)
@@
Artificial Intelligence Generated Content (AIGC) leverages artificial intelligence technologies to autonomously produce content. In this sector, we systematically classify contemporary algorithms based on the types of prompts and the objects they generate. The categories include Image-to-3D Object Generation, Text-to-3D Object Generation, Multi-Object and Scene Generation, and 4D Generation.
@@
IV-B1 Text to 3D Objects
Extensive existing research endeavors to utilize the superior generative capabilities of 2D generative models to achieve coherent 3D content creation. Benefiting from reduced dependency on extensive 3D training data, Score Distillation Sampling-based paradigms garnered significant attention in early research [99]. Some works [95, 100, 101, 102] focus on improving the framework to apply score distillation loss to 3DGS. Building upon Score Distillation Sampling (SDS), DreamGaussian [95] ensures the geometric consistency of the generated models by extracting explicit Mesh representations from the 3DGS and refines texture in the UV space to enhance the quality of the renderings. However, the mode-seeking paradigm of score distillation frequently leads to oversaturation, excessive smoothing, and lack-detail in the generated outcomes [103, 104, 105, 106, 107]. Among them, LucidDreamer [104] addresses the challenges of over-smoothing and insufficient sampling steps inherent in traditional SDS. By introducing deterministic diffusion trajectories (DDIM [108]) and interval-based score matching mechanisms Eq. 9, it achieves superior generation quality and efficiency.
To further mitigate the inherent limitations of SDS, several works [109, 110, 111, 112, 113] seek to leverage video or multi-view generative models to obtain more data for reconstruction. Although these approaches introduce direct prior guidance for 3D generation, the inherent lack of guaranteed 3D consistency in both multi-view and video generation still leads to instability in reconstruction. To enhance the efficiency of 3D asset generation, works [114, 115, 116] aim to generate using only feed-forward networks without the need for scene-specific training. BrightDreamer [114] predicts positional offsets following fixed initialization and employs a text-guided triplane generator to process extracted textual features for predicting additional 3DGS attributes, achieving text-to-3D model conversion in 77ms, albeit at the cost of some reconstruction quality. For enhanced geometric detail control, SketchDream [117] proposes a framework for sketch-based text-to-3D generation and editing, which integrates hand-drawn sketches and text prompts to achieve fine-grained geometry and appearance control, enabling high-quality 3D content creation and local editing through a two-stage coarse-to-fine approach.
More directly, several works [118, 119, 120, 121] aim to incorporate 3DGS representations into 3D generative models. Among these, L3DG [120] proposes a latent diffusion framework for compressed 3D Gaussian representation, achieving superior visual quality and real-time rendering. However, these approaches are inherently limited by the availability of 3D data, which impacts their reconstruction capability for complex targets.
@@ IV-B1 Text to 3D Objects
Some works [92, 93, 122] also attempt to apply this generative paradigm to areas such as digital human generation. HumanGaussian [93] combines RGB and depth rendering to improve the SDS, thereby jointly supervising the optimization of the structural perception of human appearance and geometry.
IV-B2 Image to 3D Object
Similar to works on NeRF, recent studies [123, 97, 124] have also focused on generating entire 3DGS from a single image. These approaches share fundamental similarities with text-to-3D object methods. As an example, following a process similar to DreamGaussian [95], Repaint123 [97] employs Zero-123 [125] and SDS for coarse 3DGS, followed by a fine stage where mesh representation is extracted and refined using depth-guided and visibility-aware repainting on novel views for consistent 3DGS fine-tuning.
IV-B3 Multi-Object and Scene Generation